2g of ammonium chloride is dissolved in 20cm3 of tap water whose initial temperature is 23°C. What is the temperature of the solution after all the ammonium chloride has dissolved?


1
Expert's answer
2021-10-17T11:25:02-0400

Molar mass of NH4Cl = 53.491

2/53.491= 0.0374mol

20/1000= 0.02

= 0.0374/0.02= 1.87

= 23×1.87= 42.998°C